OpenStudy (anonymous):

How do I simplify this? a^5(a^4)=a^[1]

OpenStudy (jhannybean):

\[a^5(a^4)=a^1\] Recall th rule for multiplying variables with the same base. \(x^n \cdot x^m = x^{n+m}\) Therefore.. \(a^5\cdot a^4 = a^{5+4} =a^9\)\[a^9=a\]\[a^9 - a = 0\] can you solve it from here?

OpenStudy (anonymous):

I think so. c: Thanks!!

OpenStudy (jhannybean):

No problem :D